# Pulse-PPG: An Open-Source Field-Trained PPG Foundation Model for Wearable Applications Across Lab and Field Settings

### v1

@@ 5.1. Downstream Datasets and Tasks
In order to evaluate the generalizability of our Pulse-PPG foundation model, we evaluate our 5 downstream datasets and 11 tasks, which can be seen in Table 1. Our datasets have a range of different PPG settings.
@@
Table 1. Downstream Datasets and Tasks. $(\boldsymbol{\cdot})$ designates either a regression task or # of classes for a classification task.
@@
Additionally, with each of PPG type, there are a set of canonical downstream tasks associated with them. For wearables, that includes stress classification (Schmidt et al., 2018a; Neupane et al., 2024), activity classification (Neupane et al., 2024; Reiss et al., 2019a), and instantaneous HR regression (Reiss et al., 2019a). For clinical settings, this includes sleep disturbance classification (Garde et al., 2014), blood pressure regression (Elgendi et al., 2022), and hypertension  (Elgendi et al., 2022). While activity classification is traditionally performed using IMU sensors or motion signals, we include it as a downstream task to showcase the versatility of our model in learning meaningful representations solely from PPG data. Each of these canonical tasks are captured in our downstream datasets. Please see Appendix A.1 for further dataset details and A.2 for evaluation implementation details.
5.1.1. Metrics
For evaluating our models across diverse downstream tasks, we utilize task-specific metrics for both regression and classification problems. For regression tasks, we report standard Mean Absolute Error (MAE) and Mean Squared Error (MSE). Additionally, we use Mean Absolute Percentage Error (MAPE) to assess relative errors across different scales.
@@ 5.1.1. Metrics
For classification tasks, we report macro F1 score, Accuracy, Precision, and Recall to evaluate overall model performance and the trade-offs between false positives and false negatives. To provide a more comprehensive assessment, particularly for imbalanced datasets, we compute Area Under the Precision-Recall Curve (AUPRC) and Area Under the Receiver Operating Characteristic Curve (AUROC), which measure the model’s ability to distinguish between classes.
5.2. Experiment 1: Pulse-PPG vs. A Prior PPG foundation model
@@ 5.2.1. Background
In order to evaluate the effectiveness of our Pulse-PPG foundation model, we benchmark against the current state-of-the-art open PPG foundation model, PaPaGei (Pillai et al., 2024). Their PaPaGei-S model uses a 1D ResNet backbone and leverages a novel self-supervised objective with PPG-morphology-specific metrics, like the Inflection Point Area ratio, achieving top performance on primarily, clinically focused downstream tasks. Crucially, they release their model weights and code to the public (https://zenodo.org/records/13983110), which we utilize for our experiments.
However, their pre-training dataset was composed solely of pre-processed, clean clinical PPG signals, and they do not demonstrate generalizability to wearable field PPG settings. As previously discussed in Section 3.2.2, we hypothesize that their PPG-morphology-specific metrics will be unable to generalize well in the much noisier, field setting. This is unlike our learned distance metric, which can potentially learn to do motif-similarity comparisons, even in noisy settings.
We also note that while PaPaGei demonstrated strong performance across various datasets in their work, their evaluation methodology involved using a different version of the foundation model for each task. Thus, while PaPaGei sets a critical stepping stone for PPG foundation modeling, our goal is to build a single open-source1 PPG foundation model that can generalize across many tasks, without re-training the network from scratch.
Figure 3. Comparison of Pulse-PPG vs. PaPaGei, another PPG foundation model (Pillai et al., 2024). The two plots shows relative performance of a linear probe evaluation for each model, with the left for classification via F1 score and the right for regression via Mean Average Percentage Error. 10/11 of the task data points reside above the slope, demonstrating how Pulse-PPG consistently outperforms PaPaGei, with particularly substantial improvements for the Wearable Field and Lab PPG tasks. See further discussion in Section 5.2.3.
@@ 5.2.2. Experimental Design
To ensure a fair comparison, we utilize a linear probing evaluation for each downstream task for both the Pulse-PPG and PaPaGei-S model. Linear probing is a widely accepted evaluation method for foundation models because it isolates the quality of each foundation model’s learned representations by freezing the model’s embeddings and training only a simple logistic classifier or linear regression. Both PaPaGei and Pulse-PPG embed PPG sequences into a single 1D 512-dimensional vector, making their linear probe results directly comparable. By using linear probing—training a simple classifier on frozen embeddings—we isolate the quality of the learned representations, avoiding biases from architectural differences.
To re-emphasize, this linear probing evaluation is not intended to achieve maximum task-specific performance, instead serving as a controlled experiment to enable a fair comparison between foundation models.
@@ 5.2.3. Results
Figure 3 visualizes the difference in performance between our Pulse-PPG foundation model versus the publicly released pre-trained PaPaGei. Full results, can be found in the 1st and 4th data columns (i.e. Wearable Field Pulse-PPG and Clinical PaPaGei) in Table 3 in the Appendix. Our Pulse-PPG achieves much stronger performance, outperforming PaPaGei on 10/11 of our downstream tasks, consistently across the many reported metrics.
@@
On the Wearable PPG tasks, our Pulse-PPG achieves a consistently large ¿.10 increase to F1 scores in the classification tasks (Stress - MOODS: 0.10 / Activity - MOODS: 0.15 / Activity - PPG-DaLiA: 0.12 / Stress (2) - WESAD: 0.15 / Stress (4) - WESAD: 0.14). Then on the instantaneous HR regression task, our Pulse-PPG model has large 10.24 improvement in MAE. These heart rate estimates are conducted on unfiltered wearable PPG data, which is significant due to the widely recognized challenge of noise-induced degradation in heart rate estimation. On the Clinical PPG tasks, our Pulse-PPG model continues to outperform PaPaGei, with $\sim 0.5$ improvement to MAE for Systolic BP, Diastolic BP, and average HR regression, as well as a .05 improvement to F1 score on hypertension classification. While clinical evaluation tasks show smaller improvements compared to wearable ones, the consistent gains highlight the model’s cross-domain generalizability.
@@
The only task on which Pulse-PPG underperforms is the Sleep Disturbance task. This can be attributed to the nature of Pulse-PPG’s pre-training dataset, MOODS. During the MOODS study, participants were instructed to wear the smartwatch during waking hours (Neupane et al., 2024). Consequently, Pulse-PPG was not exposed to PPG signals during sleep in its pre-training phase, limiting its ability to learn sleep-related patterns, since human physiology during sleep is significantly different than that of waking hours due to natural biological processes (Smets et al., 2018). In contrast, PaPaGei was pre-trained on the MESA dataset (Chen et al., 2015), which is a sleep study dataset.
5.3. Experiment 2: Assessing Field-to-Lab Generalizability
@@ 5.3.1. Background:
Figure 4. Comparison of pre-training Pulse-PPG on Wearable Field data vs. Clinical Data. The two plots shows relative performance of a linear probe evaluation for each model, with the left for classification via F1 score and the right for regression via Mean Average Percentage Error. 10/11 of the task data points reside above the slope, demonstrating how pre-training on wearable field data is highly generalizable, demonstrating strong field-to-lab generalizability and even field-to-clinical generalizability. This is contrary to the current standard line of research, which tends to focus on lab-to-field methods. See further discussion in Section 5.3.3.
@@ 5.3.2. Experimental Design
In these experiments, we assess how pre-training on wearable field PPG compares to pre-training on clean PPG. Similar to the prior experiments in Section 5.2, we utilize a linear probe for evaluation, to ensure fair comparisons. From the prior experiments, we already have our Pulse-PPG pre-trained on field PPGs and PaPaGei pre-trained on the clinical PPGs.
In order to construct a large clean clinical PPG pre-training dataset for re-training Pulse-PPG, we utilize the curated MIMIC-III PPG waveform dataset from (Xu et al., 2022). This dataset is composed of 151,738 100 Hz 5-minute-long clean clinical PPGs from 18,210 patients. [^2] Unfortunately, PaPaGei did not release their full curation code for constructing their clinical pre-training dataset, making it impossible for us to train on the same clinical data. However, MIMIC-III waveform did form a portion of their dataset. We then use this clean clinical PPG dataset to pre-train our Pulse-PPG model, following the same training procedures and hyperparameters that we used when we were pre-training on the MOODS field PPG dataset.
Next, we pre-train PaPaGei from scratch with our MOODS field dataset, using its original training procedure from their codebase located here: https://github.com/Nokia-Bell-Labs/PaPaGei-foundation-model.

### v2

@@ 5.1. Downstream Datasets and Tasks
In order to evaluate the generalizability of our Pulse-PPG foundation model, we evaluate our 5 downstream datasets and 11 tasks, which can be seen in Table 1. Our datasets have a range of different ”PPG type” settings:
@@
Table 1. Downstream Datasets and Tasks. $(\bm{\cdot})$ designates either a regression task or # of classes for a classification task.
@@
Additionally, with each of the PPG types, there is a set of canonical downstream tasks associated with them. For wearables, that includes stress classification (Schmidt et al., 2018a; Neupane et al., 2024), activity classification (Neupane et al., 2024; Reiss et al., 2019a), and instantaneous HR regression (Reiss et al., 2019a). For clinical settings, this includes sleep disturbance classification (Garde et al., 2014), blood pressure regression (Elgendi et al., 2022), and hypertension  (Elgendi et al., 2022). While activity classification is traditionally performed using IMU sensors or motion signals, we include it as a downstream task to showcase the versatility of our model in learning meaningful representations solely from PPG data. Each of these canonical tasks is captured in our downstream datasets. Please see Appendix A.1 for further dataset details.
5.2. Evaluation Implementation Details
All PPG signals were resampled to match the 50 Hz sampling frequency present in our pre-training MOODS dataset (Neupane et al., 2024). For our evaluations using PaPaGei, we made sure to resample the signals to 125 Hz to match the sampling frequency that they trained on (Pillai et al., 2024).
To train and evaluate our linear probes, we combined the train and validation splits and used a cross-validation hyperparameter grid search with the Scikit Learn package. For the regression tasks, we utilized a linear regression with an L2 regularization and conducted a grid search over the following parameters: ’alpha’:[0.1,1.0,10.0,100.0],’solver’:[’auto’,’cholesky’,’sparse_cg’] with the scoring function as negative MSE. For the classification tasks, we utilized a logistic regression with an L2 regularization and lbfgs solver and conducted a grid search over the following parameters ’C’:[0.01,0.1,1,10,100],’solver’:[’lbfgs’],’max_iter’:[1000,10_000] with the scoring function as macro F1. Inputs into the probes were normalized with the Scikit Learn StandardScaler module.
In order to train and evaluate our fine-tuning results, we fit a classification or regression head on top of the existing network, then train the entire network end-to-end. For classification, the head was made up of a layer normalization layer, followed by a linear layer, with the loss function being a linear combination of Cross Entropy loss, a soft F1 loss (Bénédict et al., 2021), and Dice loss (Li et al., 2019). For regression, the head was an MLP layer, with an initial layer with an output dimensionality of 128, followed by the GeLU activation, followed by a final linear layer. The loss function used was the MSE loss function. For both evaluations, we utilize an Adam optimizer. The checkpoint with the highest macro F1 score and the lowest MAE on the validation split was selected for inference on the test set. For exact implementation details, please refer to our public codebase: https://github.com/maxxu05/pulseppg.
5.3. Evaluation Metrics
@@
For evaluating our models across diverse downstream tasks, we utilize task-specific metrics for both regression and classification problems. For regression tasks, we report standard Mean Absolute Error (MAE) and Mean Squared Error (MSE). Additionally, we use Mean Absolute Percentage Error (MAPE) to assess relative errors across different scales. For classification tasks, we report macro F1 score, Accuracy, Precision, and Recall to evaluate overall model performance and the trade-offs between false positives and false negatives. To provide a more comprehensive assessment, particularly for imbalanced datasets, we compute Area Under the Precision-Recall Curve (AUPRC) and Area Under the Receiver Operating Characteristic Curve (AUROC), which measure the model’s ability to distinguish between classes.
6. Experimental Results
Our primary claims are that (1) Pulse-PPG is able to achieve state-of-the-art performance across a wide set of downstream PPG tasks for wearable and clinical applications, (2) pre-training on wearable field data achieves surprisingly strong performance compared to cleaner clinical data, and that (3) researchers will be able to utilize our open-sourced model in a variety of use-cases to expedite their own research efforts.
We evaluate these claims through a series of experiments: Pulse-PPG vs. PaPaGei, a prior PPG FM (Section 6.1); Pulse-PPG vs. General Time-series FMs and Supervised Baselines (Section 6.2); Assessing Field-to-Lab Generalizability (Section 6.3); and finally the Evaluation of Pulse-PPG use-cases (Section 6.4).
6.1. Experiment 1: Pulse-PPG vs. PaPaGei, A Prior PPG foundation model
@@ 6.1.1. Background
To evaluate the effectiveness of our Pulse-PPG foundation model, we benchmark against the current state-of-the-art in open PPG foundation model, PaPaGei (Pillai et al., 2024). Their PaPaGei-S model uses a 1D ResNet backbone and leverages a novel self-supervised objective with PPG-morphology specific metrics, like the Inflection Point Area ratio, achieving top performance on clinically-focused downstream tasks. Crucially, they release model weights and code (https://zenodo.org/records/13983110), which we utilize for replication.
However, their pre-training dataset was composed solely of pre-processed, clean clinical PPG signals. As previously discussed in Section 3.2.2, we hypothesize that their PPG-morphology specific metrics will be unable to generalize well in the noisier field setting. This is unlike our learned distance metric, which can potentially learn to do motif-similarity comparisons, even in noisy settings.
We also note that while PaPaGei demonstrated strong performance across various datasets in their work, their evaluation methodology involved using multiple different hyperparameter-tweaked versions of the foundation model and choosing the best model per task (Pillai et al., 2024). We report results on one set of pre-trained weights per model (for both PaPaGei and Pulse-PPG) throughout our entire paper, which explains any discrepancies in our reported numbers versus their original work. While PaPaGei sets a critical stepping stone for PPG foundation modeling, our goal is to build a single open-source2 PPG foundation model that can generalize across many tasks.
[…]
This figure compares the Pulse-PPG model with the PaPaGei foundation model, showcasing performance in classification (F1 score) and regression (MAPE). Pulse-PPG consistently outperforms PaPaGei, especially in Wearable Field and Lab PPG tasks.
Figure 3. Comparison of Pulse-PPG vs. PaPaGei (Pillai et al., 2024), a prior PPG foundation model. The two plots show the relative performance of a linear probe evaluation for each model, with the left for classification via F1 score and the right for regression via Mean Average Percentage Error. 10/11 of the task data points reside above the slope, demonstrating how Pulse-PPG consistently outperforms PaPaGei, with particularly substantial improvements for the Wearable Field and Lab PPG tasks. See further discussion in Section 6.1.3.
@@ 6.1.2. Experimental Design
To ensure a fair comparison, we utilize a linear probing evaluation for each downstream task for both Pulse-PPG and PaPaGei-S models. Linear probing is a widely accepted evaluation method for foundation models because it isolates the quality of each foundation model’s learned representations by freezing the model’s embeddings and training only a simple logistic classifier or linear regression. Both PaPaGei and Pulse-PPG embed PPG sequences into a single 1D 512-dimensional vector, making their linear probe results directly comparable. By using linear probing—training a simple classifier on frozen embeddings, we isolate the quality of the learned representations, avoiding biases from architectural differences.
To re-emphasize, this linear probing evaluation is not intended to achieve maximum task-specific performance, instead serving as a controlled experiment to enable a fair comparison between the foundation models.
@@ 6.1.3. Results
Figure 3 illustrates the difference in performance between our Pulse-PPG foundation model and the publicly released pre-trained PaPaGei. Full results can be found in Table 7 in the Appendix. Pulse-PPG outperforms PaPaGei on 10/11 of our downstream tasks, consistently across many of the reported metrics.
@@
On the Wearable PPG tasks, our Pulse-PPG achieves a consistently large .$\geq.{\color[rgb]{0,0,0}10}$ increase in F1 scores in the classification tasks, except one. Then, on the instantaneous HR regression task, our Pulse-PPG model has a large 6.28 improvement in MAE. These heart rate estimates are conducted on unfiltered wearable PPG data, which is significant due to the widely recognized challenge of noise-induced degradation in heart rate estimation. On the Clinical PPG tasks, our Pulse-PPG model continues to outperform PaPaGei, with $\sim 0.5$ improvement in MAE for Systolic BP, Diastolic BP, and average HR regression, as well as a .05 improvement in F1 score on hypertension classification. While clinical evaluation tasks show smaller improvements compared to wearable ones, the consistent gains highlight the model’s cross-domain generalizability.
@@
The only task on which Pulse-PPG underperforms is the Sleep Disturbance task. This can be attributed to the nature of Pulse-PPG’s pre-training dataset, MOODS, where participants were instructed to wear the smartwatch only during waking hours (Neupane et al., 2024). Consequently, Pulse-PPG was not exposed to PPG signals during sleep in its pre-training phase, and human physiology during sleep is significantly different from that of waking hours due to natural biological processes (Smets et al., 2018). In contrast, PaPaGei’s pre-training included a sleep study dataset (Chen et al., 2015).
6.2. Experiment 2: Pulse-PPG vs. General Time-series FMs and Supervised Baselines
Table 2. Pulse-PPG vs. PPG FM, General Time-series FM, and Supervised Baselines. Mean Performance across Classification and Regression Tasks. The best is bolded, and the second is underlined. Full results can be found in Table  6. Pulse-PPG achieves stronger performance across all metrics than the prior PPG foundation model, PaPaGei. It always maintains Top-2 performance, even among general time-series foundation models, which have significantly higher parameter count (Pulse-PPG with 28.5m vs. Chronos with 200m and MOMENT with 385m) and among fully-supervised models.
Avg. Classification Performance Avg. Regression Performance
Model F1 Score Accuracy Precision Recall AUPRC AUROC MAE MSE MAPE
PPG FM Pulse-PPG $\bm{{\color[rgb]{0,0,0}0.5870}\pm 0.17}$ $\bm{{\color[rgb]{0,0,0}0.6840}\pm 0.17}$ ${\color[rgb]{0,0,0}0.6237}\pm 0.16$ $\bm{{\color[rgb]{0,0,0}0.5894}\pm 0.17}$ ${\color[rgb]{0,0,0}0.5149}\pm 0.20$ $\bm{{\color[rgb]{0,0,0}0.7481}\pm 0.17}$ ${\color[rgb]{0,0,0}8.8593}\pm 6.25$ ${\color[rgb]{0,0,0}145.255}\pm 171.08$ ${\color[rgb]{0,0,0}0.1009}\pm 0.05$
PaPaGei(Pillai et al., 2024) $0.4994\pm 0.14$ $0.6361\pm 0.15$ $0.5530\pm 0.14$ $0.5210\pm 0.14$ $0.4164\pm 0.12$ $0.6840\pm 0.10$ $10.7488\pm 7.75$ $204.6825\pm 237.78$ $0.1288\pm 0.10$
General FM Chronos(Ansari et al., 2024) ${\color[rgb]{0,0,0}0.5845}\pm 0.12$ ${\color[rgb]{0,0,0}0.6755}\pm 0.14$ $\bm{{\color[rgb]{0,0,0}0.6386}\pm 0.14}$ ${\color[rgb]{0,0,0}0.5875}\pm 0.13$ $\bm{{\color[rgb]{0,0,0}0.5189}\pm 0.16}$ ${\color[rgb]{0,0,0}0.7329}\pm 0.15$ ${\color[rgb]{0,0,0}9.1902}\pm 6.40$ ${\color[rgb]{0,0,0}162.1719}\pm 211.33$ ${\color[rgb]{0,0,0}0.1054}\pm 0.04$
MOMENT(Goswami et al., 2024) ${\color[rgb]{0,0,0}0.5456}\pm 0.15$ ${\color[rgb]{0,0,0}0.6691}\pm 0.15$ ${\color[rgb]{0,0,0}0.6059}\pm 0.15$ ${\color[rgb]{0,0,0}0.5603}\pm 0.15$ ${\color[rgb]{0,0,0}0.4902}\pm 0.19$ ${\color[rgb]{0,0,0}0.7241}\pm 0.16$ $\bm{{\color[rgb]{0,0,0}8.5742}\pm 6.09}$ $\bm{{\color[rgb]{0,0,0}139.6036}\pm 168.57}$ $\bm{{\color[rgb]{0,0,0}0.0988}\pm 0.05}$
Supervised 1D ResNet-26 ${\color[rgb]{0,0,0}0.4623}\pm 0.16$ ${\color[rgb]{0,0,0}0.6212}\pm 0.17$ ${\color[rgb]{0,0,0}0.4684}\pm 0.18$ ${\color[rgb]{0,0,0}0.4944}\pm 0.14$ ${\color[rgb]{0,0,0}0.4025}\pm 0.15$ ${\color[rgb]{0,0,0}0.6224}\pm 0.18$ ${\color[rgb]{0,0,0}17.2733}\pm 18.49$ ${\color[rgb]{0,0,0}580.3867}\pm 961.79$ ${\color[rgb]{0,0,0}0.1803}\pm 0.15$
Rand Forest ${\color[rgb]{0,0,0}0.4152}\pm 0.10$ ${\color[rgb]{0,0,0}0.5911}\pm 0.15$ ${\color[rgb]{0,0,0}0.4671}\pm 0.12$ ${\color[rgb]{0,0,0}0.4337}\pm 0.10$ ${\color[rgb]{0,0,0}0.3328}\pm 0.08$ ${\color[rgb]{0,0,0}0.5977}\pm 0.09$ ${\color[rgb]{0,0,0}12.5655}\pm 7.70$ ${\color[rgb]{0,0,0}268.2220}\pm 283.46$ ${\color[rgb]{0,0,0}0.1495}\pm 0.08$
6.2.1. Background
To evaluate the effectiveness of Pulse-PPG, we compare it against additional baseline models. We include general-purpose time series foundation models, which are widely adopted for benchmarking due to their large-scale and diverse pretraining datasets, in this comparison. Specifically, we benchmark against Chronos (Ansari et al., 2024) and MOMENT (Goswami et al., 2024), two well-known general-purpose time series foundation models in this space. Chronos trains language models with suitable adaptations to time series data from domains that include energy, finance, economics, nature, retail, and transportation. MOMENT uses reconstruction of masked embeddings as a pre-training task, on datasets covering healthcare, human physiology, nature, power, economics, traffic, and weather. For supervised baselines, from classical machine learning, we use Random Forest, while from deep learning, we employ 1D ResNet-26, the same backbone used in the original Pulse-PPG model.
6.2.2. Experimental Design:
We train a Random Forest model with statistical features—mean, median, maximum, minimum, 25th percentile, and 75th percentile, computed from raw PPG windows corresponding to ground truth from respective downstream datasets. To train the ResNet model, we append a dense layer with output units equal to the number of classes, followed by an activation layer. For regression tasks, we use a dense layer with a single output unit without activation. Chronos (200 million) and MOMENT (385 million) produce embeddings of size 768 and 1024, respectively (compared to Pulse-PPG’s 28.5 million parameters with an embedding size of 512). We extract embeddings from these models for respective input windows from each of the downstream datasets and apply the same linear probe pipeline for classification and regression tasks.
6.2.3. Results:
We present downstream task performance of Pulse-PPG against supervised baselines and general-purpose time series models in Table 6 in the Appendix. We summarize the results with their average task-relevant (classification/regression) performance metrics in Table 2. Pulse-PPG was trained on a single dataset in a self-supervised fashion, yet it demonstrates balanced generalizability across diverse downstream tasks with a simple linear probe evaluation. Whereas, 1D ResNet-26 and Random Forest were trained separately on each downstream dataset. While ResNet-26 achieves exceptional performance for certain tasks, e.g., an F1-score of 0.7903 for PPG-DaLiA multiclass activity detection, its performance in multiclass stress classification for WESAD with an F1-score of 0.1998 is much worse. Pulse-PPG outperforms the Random Forest model on the majority of tasks.
General-purpose time series models benefit from exposure to diverse datasets, both in terms of nature and scale. This diversity enables them to learn periodicities and trends across contexts, helping them capture complex temporal patterns. Hence, although general-purpose time series models may outperform modality-specific foundation models like Pulse-PPG and PaPaGei in certain tasks, e.g., multi-class activity detection, the value of pretraining on noisy, real-world field PPG data is evident. As shown in Table 2, Pulse-PPG consistently ranks among the top two across a wide range of tasks focused on diverse applications in the physiological domain.
6.3. Experiment 3: Assessing Field-to-Lab Generalizability
@@ 6.3.2. Experimental Design
In these experiments, we assess how pre-training on wearable field PPG compares to pre-training on clean PPG. Similar to the prior experiments in Section 6.1, we utilize a linear probe for evaluation to ensure fair comparisons. From the prior experiments, we already have our Pulse-PPG pre-trained on field PPGs and PaPaGei pre-trained on the clinical PPGs.
In order to construct a large clean clinical PPG pre-training dataset for re-training Pulse-PPG, we utilize the curated MIMIC-III PPG waveform dataset from (Xu et al., 2022). This dataset is composed of 151,738 100 Hz 5-minute-long clean clinical PPGs from 18,210 patients. [^4] Unfortunately, PaPaGei did not release their full curation code for constructing their clinical pre-training dataset, making it impossible for us to train on the same clinical data. However, the MIMIC-III waveform did form a significant portion of their dataset. We then use this clean clinical PPG dataset to pre-train our Pulse-PPG model, following the same training procedures and hyperparameters that we used when we were pre-training on the MOODS field PPG dataset.
Next, we pre-train PaPaGei from scratch with the MOODS field dataset, using its original training procedure from their codebase located here: https://github.com/Nokia-Bell-Labs/PaPaGei-foundation-model.
[…]
This figure compares the generalizability of models pre-trained on wearable field data versus clinical data. It shows classification performance via F1 score and regression accuracy via MAPE, highlighting better field-to-lab generalizability from wearable data.
Figure 4. Comparison of Pre-training on Wearable Field data vs. Clinical Data to Assess Field-to-Lab Generalizability. The two plots show the relative performance of a linear probe evaluation for each model, with the left for classification via F1 score and the right for regression via Mean Average Percentage Error. 10/11 of the task data points reside above the slope, demonstrating how pre-training on wearable field data is highly generalizable, demonstrating strong field-to-lab generalizability and even field-to-clinical generalizability. This is contrary to the current standard line of research, which tends to focus on lab-to-field methods. See further discussion in Section 6.3.3.